ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 80A30, 82C41 Fedotov, Sergei, Mathematics Department, UMIST, Manchester, UK Travelling Waves in a Reaction-Diffusion System with KPPkinetics A new asymptotic method is presented for the analysis of the travelling waves in the reaction-diffusion system with the diffusion with a finite velocity and Kolmogorov-Petrovskii-Piskunov kinetics. The analysis makes use of the functional integral approach, scaling procedure and the singular perturbation techniques involving the large deviations theory for the Poisson random walk. The exact formula for the position and speed the reaction front is derived. It is found that the reaction front dynamics is formally associated with the relativistic Hamiltonian mechanics.